# vTrain: A Simulation Framework for Evaluating Cost-effective and Compute-optimal Large Language Model Training

### v1

@@ V-B Case Study #2: Cost-effective Multi-tenant GPU Clusters
We employ a simulation-based methodology to demonstrate vTrain’s effectiveness because the estimated cost of running the same experiments we present in this section at Amazon AWS is over 100 million dollars (we study a total of $41$ workload traces, each trace modeling a GPU cluster operating for 400 hours of wall clock training time using 1,024 A100 GPUs) [^1] It is worth emphasizing that ElasticFlow, due to similar cost reasons, demonstrated their proposal’s efficacy using simulation. We assume a GPU cluster with 128 nodes (each node containing $8$ A100 GPUs, identical to our multi-node validation study in Section IV), amounting to 1,024 GPUs in total. Our simulator evaluates the entire lifetime of a training job, from its arrival to its completion. In terms of the workloads we evaluate, we use real-world traces from Microsoft’s internal ITP clusters [31] to model the arrival time of each training job. Specifically, a single workload trace is generated by randomly sampling $N$ consecutive job arrival points from Microsoft’s trace which are used to model the time intervals in which two consecutive jobs arrive to the training cluster. All workload traces are assumed to arrive within a fixed time period, meaning a workload traces with larger number of jobs will likely incur larger deadline violation rates (i.e., workload trace with $128$ jobs will stress the training cluster more than $64$ jobs thereby invoking larger number of deadline violations). Each arrived job is randomly chosen among the three LLM configurations listed in Table II whose total number of training iterations to execute and its requested deadline guarantee is also randomly chosen.
TABLE II: Model configurations and batch sizes used in our GPU cluster experiments in Section V-B. $L$ is the number of transformer layers, $h$ is the size of hidden dimensions, $n$ is the number of attention heads, $s$ is the maximum sequence length, and $B$ is the batch size used for the LLM training job.
Parameters
$L$ $h$ $n$ $s$ $B$
@@
We implement the exact same scheduling algorithm ElasticFlow proposes, which elastically adjusts the GPU resources allocated to each training job based on the profiled training performance numbers. The difference in how well baseline ElasticFlow vs. vTrain optimize scheduling decisions primarily lies in how close the best profiled training performance is to the performance achievable with an optimal parallelization plan. For baseline ElasticFlow, limiting its profiled parallelization plans only to the data parallelism dimension while fixing the tensor/pipeline parallelism dimension value to “one” (which is what ElasticFlow assumes in [15]) severely constrains the possible design points the GPU cluster scheduler can leverage for scheduling decisions. We therefore have baseline ElasticFlow keep a minimum tensor/pipeline parallelism degree necessary for each LLM model configuration, e.g., the baseline ElasticFlow policy parallelizes the $39.1$B LLM model in Table II across ($16$$\times$$d$) GPUs using 8-way tensor parallelism, 2-way pipeline parallelism, and $d$-way data parallelism. Our vTrain, on the other hand, can utilize its knowledge of the optimal parallelization plan for scheduling, guaranteeing at a minimum to provide the same training performance that baseline ElasticFlow can provide. Below we detail the effectiveness of vTrain in terms of deadline satisfactory rate, job completion time, and makespan.
@@
[…]
Job completion time (JCT) of a job is the time spent from its arrival to its completion. Since ElasticFlow simply terminates jobs that cannot be finished within their deadlines (which can artificially lower JCT), deriving JCT under a deadline-free setting is another important metric to measure for a fair comparison. Thus, we generate nine 32-job traces (without specifying job deadlines) which has a relatively lenient traffic condition. As shown in Figure 13, vTrain reduces JCT by an average 52.59% while always guaranteeing JCT to become shorter than ElasticFlow.
Makespan is the time spent until all jobs are completely executed. For the purpose of this study, we generate five synthetic workload traces (containing $16$ to $72$ jobs) where the $N$ jobs are all submitted to the GPU cluster at the same time. We then measure the makespan for both systems, the result of which is summarized in Figure 14. For the 16-job workload which represents the lightest traffic, ElasticFlow shows the least difference in makespan vs. vTrain. As the number of jobs increases, however, baseline suffers from effectively allocating limited GPU resources to the jobs due to its sub-optimal choice of parallelization plans. Our vTrain helps the cluster scheduling algorithm better utilize the scarce GPU resources and reduces makespan up to 26.43%.
@@ V-C Case Study #3: Compute-optimal LLM Model Size
Naïvely assuming $100\%$ GPU utility leads to an (overprovisioned) compute budget of $C=2.72e\text{+}24$ FLOPs over the period of 30 days. With such compute budget $C$ in place, one would lead to the conclusion that a compute-optimal model size of $N$ $=$ $145.61$ billion parameters can be trained with $T$ $=$ 2,912 billion tokens within 30 days. However, vTrain reveals that, even with the most optimal parallelization strategy, the training system in aggregate achieves an average $30.37\%$ GPU utilization. This leads to a total of $95$ days of training time to train $T$ tokens over the LLM with $N$ parameters, more than 3×3$\times$3 × the originally expected $30$ days of training time.
TABLE III: Compute-optimal Chinchilla points under different compute budget. The “Optimal (t,d,p)” column represents the optimal 3D parallelism plan under the given compute budget and LLM model architecture.
@@
12,288 80 145.61 2,912 (8,42,10) 95
12,288 70 127.49 2,550 (8,60,7) 72
12,288 60 109.37 2,187 (8,70,6) 53
10,240 80 101.21 2,024 (8,42,10) 50
10,240 70 88.62 1,772 (8,84,5) 37
10,240 60 76.04 1,521 (8,105,4) 27
9,216 80 82.03 1,641 (8,84,5) 34
9,216 70 71.83 1,437 (8,84,5) 26
9,216 60 61.64 1,233 (8,105,4) 19
8,192 70 56.81 1,136 (8,84,5) 18
8,192 60 48.75 975 (70,6,1) 13
Realistically estimating compute-optimal Chinchilla point. We now discuss how vTrain can be utilized to evaluate a more realistic value of $N$ and $T$ by taking effective GPU utilization into consideration. We begin by exploring different model sizes, varying the number of decoder layers ($L$) and hidden dimensions ($h$) of LLMs. For a given LLM model configuration, we then exhaustively evaluate all possible parallelism strategies within the ($t$, $d$, $p$)-way 3D parallelism search space. By identifying the parallelism strategy that results in the smallest single-iteration training time, we check whether the end-to-end training time to consume the entire $T$ tokens falls within the “effective” compute budget $C$. Among all candidate $N$, $T$, and $C$ Chinchilla points, the design point with the largest model size $N$ becomes the most compute-optimal LLM with the highest algorithmic performance given the compute budget $C$. Table III summarizes such exploration results where the first row with the estimated training time of $95$ days is the Chinchilla point derived assuming a naïve $100\%$ GPU compute utility. Using vTrain, we arrive at a more realistic Chinchilla point having a model size of $76.04$ billion parameters trained with $1,521$ billion tokens within $30$ days, which is a $48\%$ smaller LLM model architecture compared to the naïvely estimated version.
@@ VI Related Work
Distributed training. Several studies[27, 23, 62, 11, 54, 59, 55] have concentrated on partitioning training workloads across multiple workers. Data parallelism [26] is the most popular approach as it can be applied independently of the model architecture. GPipe [20] and PipeDream [33, 34] propose layer-wise partitioning of model parameters and scheduling approaches to effectively execute partitioned layers in a pipelined manner. Megatron-LM [60] proposed tensor parallelism which is tailored for transformer-based language models. There are also prior works proposing advanced collective communication algorithms for distributed training. FlexReduce [24], BlueConnect [6], and Themis [57] proposed novel bandwidth-aware All-Reduce algorithms on heterogeneous/asymmetric network topology. EFLOPS [10] proposed an algorithm-architecture co-designed for distributed training platforms with a topology-aware All-Reduce algorithm. Putting these efforts altogether, Microsoft and NVIDIA discussed their distributed training cluster for Megatron-Turing NLG 530B [61] and the 1T-parameter GPT model [36] on thousands of GPUs using 3D parallelism.
Simulation methodology for distributed training. While not targeting LLMs, ASTRA-sim [56] presents a cycle-level simulator targeting distributed training. Implemented on top of Garnet [1], ASTRA-sim focuses on simulating collective communication algorithms involved in different parallelization strategies. There are also several studies focusing on automatically searching efficient parallel training plans for modern AI models [67, 72] but these works either target small scale training systems with only tens of GPUs. vTrain is specifically designed to evaluate modern LLMs requiring several thousands of GPUs for training, providing a fast yet highly accurate simulation toolchain for AI practitioners.
@@
Exploiting deterministic dataflow of ML. There has been several prior work employing a profiling-based approach to exploit the deterministic execution property of modern AI algorithms for system optimizations. DeepPool [47] employs a profiling based scheduling algorithm to dynamically adjust GPU resource allocations for multi-tenant training jobs. PREMA [8] and LazyBatching [7] utilize the deterministic latency of AI inference to better schedule and batch input queries in inference servers. Daydream [73] proposes a simulation methodology that estimates the efficacy of a diverse set DNN optimizations (e.g., CPU-GPU tensor offloading, mixed precision training, etc). None of these prior works target modern, scale-out LLM training systems let alone providing an accurate prediction mechanism to evaluate cost-effective and compute-optimal LLM training plans.
@@ VII Conclusion
This work presents vTrain, a high-performance profiling-driven simulation framework that accurately estimates LLM’s training time. vTrain targets AI systems equipped with thousands of GPUs for training LLM model architectures containing billions of parameters trained with trillions of tokens. We demonstrate vTrain’s merits and its practicality by conducting several case studies, e.g., cost-effective evaluation of optimal training parallelization strategies that balances training time and its associated cost, efficient multi-tenant GPU cluster scheduling algorithms targeting multiple LLM training jobs, and determining a cost-optimal LLM model architecture given a fixed compute budget.

### v2

@@ V-B Case Study #2: Cost-effective Multi-tenant GPU Clusters
We employ a simulation-based methodology to demonstrate vTrain’s effectiveness because the estimated cost of running the same experiments we present in this section at Amazon AWS is over 100 million dollars (we study a total of $41$ workload traces, each trace modeling a GPU cluster operating for 400 hours of wall clock training time using 1,024 A100 GPUs)  [^2] It is worth emphasizing that ElasticFlow, due to similar cost reasons, demonstrated their proposal’s efficacy using simulation.. We assume a GPU cluster with 128 nodes (each node containing $8$ A100 GPUs, identical to our multi-node validation study in Section IV), amounting to 1,024 GPUs in total. Our simulator evaluates the entire lifetime of a training job, from its arrival to its completion. In terms of the workloads we evaluate, we use real-world traces from Microsoft’s internal ITP clusters [34] to model the arrival time of each training job. Specifically, a single workload trace is generated by randomly sampling $N$ consecutive job arrival points from Microsoft’s trace which are used to model the time intervals in which two consecutive jobs arrive to the training cluster. All workload traces are assumed to arrive within a fixed time period, meaning a workload traces with larger number of jobs will likely incur larger deadline violation rates (i.e., workload trace with $128$ jobs will stress the training cluster more than $64$ jobs thereby invoking larger number of deadline violations). Each arrived job is randomly chosen among the three LLM configurations listed in Table III whose total number of training iterations to execute and its requested deadline guarantee is also randomly chosen.
TABLE III: Model configurations and batch sizes used in our GPU cluster experiments in Section V-B. $L$ is the number of transformer layers, $h$ is the size of hidden dimensions, $n$ is the number of attention heads, $s$ is the maximum sequence length, and $B$ is the batch size used for the LLM training job.
Parameters (billion) $L$ $h$ $n$ $s$ $B$
@@
We implement the exact same scheduling algorithm ElasticFlow proposes, which elastically adjusts the GPU resources allocated to each training job based on the profiled training performance numbers. The difference in how well baseline ElasticFlow vs. vTrain optimize scheduling decisions primarily lies in how close the best profiled training performance is to the performance achievable with an optimal parallelization plan. For baseline ElasticFlow, limiting its profiled parallelization plans only to the data parallelism dimension while fixing the tensor/pipeline parallelism dimension value to “one” (which is what ElasticFlow assumes in [16]) severely constrains the possible design points the GPU cluster scheduler can leverage for scheduling decisions. We therefore have baseline ElasticFlow keep a minimum tensor/pipeline parallelism degree necessary for each LLM model configuration, e.g., the baseline ElasticFlow policy parallelizes the $39.1$B LLM model in Table III across ($16$$\times$$d$) GPUs using 8-way tensor parallelism, 2-way pipeline parallelism, and $d$-way data parallelism. Our vTrain, on the other hand, can utilize its knowledge of the optimal parallelization plan for scheduling, guaranteeing at a minimum to provide the same training performance that baseline ElasticFlow can provide. Below we detail the effectiveness of vTrain in terms of deadline satisfactory rate, job completion time, and makespan.
@@
Deadline satisfactory ratio is defined as the ratio of jobs that have met their deadlines, one which is used as the most important metric to highlight ElasticFlow’s deadline-aware scheduling algorithm. To evaluate this metric, we set the deadline for each job as $\lambda\cdot\textit{duration}$ after its arrival, where $\lambda$ is sampled from a uniform distribution [0.5, 1.5]. We simulate 64-job and 128-job workloads synthesized from each trace for both baseline and vTrain-enabled systems and measure the deadline satisfactory ratio. As shown in Figure 12, the vTrain-enabled system always satisfies more job requests than the baseline ElasticFlow, improving the deadline satisfactory ratio compared to the baseline system by 1.09$\times$ and 1.23$\times$ on average for 64-job and 128-job workloads, respectively.
Job completion time (JCT) of a job is the time spent from its arrival to its completion. Since ElasticFlow simply terminates jobs that cannot be finished within their deadlines (which can artificially lower JCT), deriving JCT under a deadline-free setting is another important metric to measure for a fair comparison. Thus, we generate nine 32-job traces (without specifying job deadlines) which has a relatively lenient traffic condition. As shown in Figure 13, vTrain reduces JCT by an average 15.21% while always guaranteeing JCT to become shorter than ElasticFlow.
@@
Makespan is the time spent until all jobs are completely executed. For the purpose of this study, we generate five synthetic workload traces (containing $16$ to $72$ jobs) where the $N$ jobs are all submitted to the GPU cluster at the same time. We then measure the makespan for both systems, the result of which is summarized in Figure 14. For the 16-job workload which represents the lightest traffic, ElasticFlow shows the least difference in makespan vs. vTrain. As the number of jobs increases, however, baseline suffers from effectively allocating limited GPU resources to the jobs due to its sub-optimal choice of parallelization plans. Our vTrain helps the cluster scheduling algorithm better utilize the scarce GPU resources and reduces makespan up to 23.03%.
@@ V-C Case Study #3: Compute-optimal LLM Model Size
Naïvely assuming $100\%$ GPU utility leads to an (overprovisioned) compute budget of $C=2.72e\text{+}24$ FLOPs over the period of 30 days. With such compute budget $C$, one would lead to the conclusion that a compute-optimal model size of $N$ $=$ $145.61$ billion parameters can be trained with $T$ $=$ 2,912 billion tokens within 30 days. However, vTrain reveals that, even with the most optimal parallelization, the training system in aggregate achieves an average $35.56\%$ GPU utility. This leads to a total of $85$ days of training time to train $T$ tokens over the LLM with $N$ parameters, more than 3×3$\times$3 × the originally expected $30$ days of training time.
TABLE IV: Compute-optimal Chinchilla points under different compute budgets. The “Optimal (t,d,p)” column represents the optimal 3D parallelism plan under the given compute budget and LLM model architecture.
@@
12,288 80 145.61 2,912 (8,42,10) 85
12,288 70 127.49 2,550 (8,60,7) 64
12,288 60 109.37 2,187 (8,70,6) 47
10,240 70 88.62 1,772 (8,84,5) 40
10,240 60 76.04 1,521 (8,70,6) 30
9,216 80 82.03 1,641 (8,84,5) 37
9,216 70 71.83 1,437 (8,84,5) 29
Realistically estimating compute-optimal Chinchilla point. We now discuss how vTrain can be utilized to evaluate a more realistic value of $N$ and $T$ by taking effective GPU utilization into consideration. We begin by exploring different model sizes, varying the number of decoder layers ($L$) and hidden dimensions ($h$) of LLMs. For a given LLM model configuration, we then exhaustively evaluate all possible parallelism strategies within the ($t$, $d$, $p$)-way 3D parallelism search space. By identifying the parallelism strategy that results in the smallest single-iteration training time, we check whether the end-to-end training time to consume the entire $T$ tokens falls within the “effective” compute budget $C$. Among all candidate $N$, $T$, and $C$ Chinchilla points, the design point with the largest model size $N$ becomes the most compute-optimal LLM with the highest algorithmic performance given the compute budget $C$. Table IV summarizes such exploration results where the first row with the estimated training time of $85$ days is the Chinchilla point derived assuming a naïve $100\%$ GPU compute utility. Using vTrain, we arrive at a more realistic Chinchilla point having a model size of $76.04$ billion parameters trained with $1,521$ billion tokens within $30$ days, which is a $48\%$ smaller LLM model architecture compared to the naïvely estimated version.
@@ VI Related Work
Distributed training. Several studies [29, 25, 68, 12, 60, 65, 61] explored distributed training across multiple workers. GPipe [21] and PipeDream [37, 38] propose layer-wise partitioning of model parameters and scheduling approaches to effectively execute partitioned layers in a pipelined manner. Megatron-LM [66] proposed tensor parallelism which is tailored for transformer-based language models. There are also prior works proposing advanced collective communication algorithms for distributed training. FlexReduce [26], BlueConnect [6], and Themis [63] proposed novel bandwidth-aware All-Reduce algorithms on heterogeneous/asymmetric network topology. Putting these efforts together, Microsoft/NVIDIA discussed their distributed training for Megatron-Turing NLG 530B [67] and the 1T-parameter GPT model [40] on thousands of GPUs via 3D parallelism.
Performance modeling for distributed training. While not targeting LLMs, ASTRA-sim [62] presents a cycle-level simulator targeting distributed training, focusing on modeling collective communication algorithms involved in different multi-dimensional network topologies. Instead of using the cycle-level network backend (i.e., Garnet [1]), ASTRA-sim2.0 [75] employs an analytical model for network backend, reducing simulation time. There are also several studies focusing on accurately modeling performance for modern AI models, but these works often limit their scope by either designing analytical models [36, 53, 80, 23], profiling a representative subset of operators or specific configurations to project the performance of others [53, 54, 80], or targeting small scale training systems with only tens of GPUs [74, 80, 54]. Specifically, AMPeD [36] is an analytical model for estimating performance in distributed training of transformer-based models, but its performance modeling is controlled by a factor for compute core efficiency, which is driven by fitting empirical experimental results on various transformer-based models and underlying hardwares, sacrificing specificity for individual scenarios. Conversely, Tale of Two Cs [53] employs a focused methodology to understand computation and communication scaling in future transformers, selecting representative operators for detailed profiling and using these to model performance of the entire training process across various model hyperparameters. Calculon [23] proposes a parameterized analytical performance model exploring optimal parallel training configurations. This work parameterizes the target LLM, system, and training configuration to produce results broken down into compute time and memory usage. Although validation was performed on a real multi-node GPU system, the validation was conducted over 8 data points only, and the outcomes from the suggested optimal search strategy have not been validated over real GPU systems. In contrast, vTrain and its practicality have been validated over various multi-node training systems as shown in Section IV and Section V-A (Table II). More critically, Calculon’s analytical model does not consider framework-level/low-level implementations of the target LLM’s execution because it is based on analytical modeling of a very specific implementation of LLM execution engines (e.g., Megatron), unlike vTrain where any changes or software updates to the framework-level/low-level implementations of LLMs (e.g., an upgrade from FlashAttention [11] to FlashAttention-2 [10] to accelerate attention layers) will naturally be captured by vTrain as our work employs profiling for performance estimation.
TABLE V: Comparison of vTrain vs. other performance models.
ASTRA-
sim [62]
AMPeD
[36]
SeqPoint
[54]
Tale of
Two Cs [53]
Calculon
[23]
vTrain
Target workload Any Transformer RNN Transformer Transformer Transformer
Simulation time
for a single
training iteration
N/A seconds N/A N/A milliseconds seconds
Performance
modeling
cycle-
level
(analytical
for 2.0)
analytical
profile-
based
profile-
based
analytical
profile-
based
Applicable to
different models
O X X X X O
Multi-GPU
support
O O X O O O
Validated over
hundreds of GPUs
X O X X O O
Number of
validation points
0
12 (single)
9 (multi)
18 0 8 (multi)
1,440 (single)
112 (multi)
Average
validation error
N/A $\sim$121.50% N/A 3.65%
8.37% (single)
14.73% (multi)
SeqPoint [54] proposes performance modeling for training of sequence-based DNNs, selecting and profiling representative set of training iterations for the entire training process based on the key observation that hardware behavior differs according to the input sequence length. While these works estimate the costs of un-profiled operators and configurations based on the profiling results of representative groups, vTrain basically profiles all operators, including even short-living element-wise operations, but minimizes the profiling overhead by identifying repetitive operators that are inherently present in transformer-based LLM training algorithms. In short, vTrain is specifically designed to evaluate modern LLMs requiring several thousands of GPUs for training, providing a fast yet highly accurate simulation toolchain for AI practitioners, one which we demonstrate using our case studies. Table V highlights key differences between vTrain and other performance models for AI systems.
@@ X O X X O O
Exploiting deterministic dataflow of ML. There has been prior work employing a profiling-based approach to exploit the deterministic execution property of modern AI algorithms for system optimizations. Rammer [30] proposes a DNN compiler design that leverages deterministic characteristics of operators in ML models to generate an efficient operator-level schedule at compile time, maximizing hardware utilization. DeepPool [51] employs a profiling based scheduling to dynamically adjust GPU resource allocations for multi-tenant training jobs. PREMA [8] and LazyBatching [7] utilize the deterministic latency of AI inference to better schedule and batch input queries in inference servers. Daydream [81] proposes a simulation methodology that estimates the efficacy of a diverse set DNN optimizations (e.g., CPU-GPU tensor offloading, mixed precision training, etc). None of these prior works target modern, scale-out LLM training systems nor provide an accurate prediction mechanism to evaluate cost-effective, compute-optimal LLM training plans.
@@ VII Conclusion
vTrain is a high-performance profiling-driven simulation framework that accurately estimates LLM’s training time. vTrain targets AI systems equipped with thousands of GPUs for training LLM models containing billions of parameters trained with trillions of tokens. We demonstrate vTrain’s merits and its practicality by conducting several case studies, e.g., evaluation of optimal training parallelization plans that balances training time and its associated cost, efficient multi-tenant GPU cluster scheduling algorithms targeting multiple LLM training jobs, and determining a cost-optimal LLM model architecture given a fixed compute budget.
Acknowledgment
This work was partly supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (NRF-2021R1A2C2091753), by Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.RS-2024-00402898, Simulation-based High-speed/High-Accuracy Data Center Workload/System Analysis Platform), by Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.RS-2024-00395134, DPU-Centric Datacenter Architecture for Next-Generation AI Devices), by Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.RS-2024-00438851, (SW Starlab) High-performance Privacy-preserving Machine Learning System and System Software), and by the Samsung Electronics. We also appreciate the support from Samsung Advanced Institute of Technology, Samsung Electronics Co., Ltd. Minsoo Rhu is the corresponding author.